OpenStudy (anonymous):

Find all real solutions for the following: (x+3)^2(x-1)+(x+3)(x-1)^2=0

OpenStudy (anonymous):

factoring is your friend

OpenStudy (anonymous):

\[(x+3)^2(x-1)+(x+3)(x-1)^2=0\]

OpenStudy (zzr0ck3r):

don't divide them

OpenStudy (zzr0ck3r):

(x+3) 2 (x−1)+(x+3)(x−1) 2 = (x+3)(x-1)((x+3)^2+(x-1))

OpenStudy (anonymous):

^Thanks. I was unsure of how I should go about factoring it.

OpenStudy (anonymous):

each term has a common factor of \((x+3)(x-1)\) factor it out

OpenStudy (anonymous):

you get \[(x+3)(x-1)(x-3+x+1)\] \[(x+2)(x-1)(2x-2)\]

OpenStudy (anonymous):

Like @funinabox said, factoring is your friend. Factor out (x+3) and (x-1):\[\bf (x+3)^2(x-1)+(x+3)(x-1)^2=0 \rightarrow (x+3)(x-1)[(x+3)+(x-1)]\]Simplifying yields:\[\bf (x+3)(x-1)(2x+2)=2(x+3)(x-1)(x+1)\]

OpenStudy (zzr0ck3r):

(x+3)^2(x-1)+(x+3)(x-1)^2=0 @satellite73 (x+3)(x−1) ((x+3)^2+x-1)

OpenStudy (anonymous):

yeah i screwe, up

OpenStudy (anonymous):

@genius12 has it

OpenStudy (anonymous):

Thanks everyone!

OpenStudy (anonymous):

should be \[(x+2)(x-1)(2x+2)\]

OpenStudy (anonymous):

@satellite73 u screwed up again lol.